### v2

@@ 5.3 Experimental Setup
Experiments are conducted on NVIDIA RTX 4090 GPUs. Specifically, we use a single GPU for image classification and crowd counting tasks, while object detection is trained in parallel using 8 RTX 4090 GPUs. For the image classification task, we adopt the Vision Transformer (ViT) [4] as our baseline method. Across all experiments, we set the learning rate to $1\times 10^{-4}$ and use the Adam optimizer [56] with a cosine annealing learning rate scheduler. Standard data augmentation techniques, including random cropping, random scaling, and horizontal flipping, are applied during training. For the crowd counting task, we adopt CCTrans [39] as our baseline method, which is a representative Transformer-based crowd counting framework. We strictly follow the original experimental setup of CCTrans to ensure fairness in comparison. For the object detection task, we select state-of-the-art detection models YOLO11 [57] and YOLOv12 [58] as baselines. We faithfully reproduce and evaluate these models under their original experimental settings [57, 58] to ensure a fair and rigorous comparison. When testing the latency and FLOPs of the model, we utilized one Tesla T4 GPU with TensorRT FP16, also the same as YOLO11 and YOLOv12 official settings.
By default, both the number of heads and the number of soft hyperedges in SoftHGNN are set to 8. In the variant with Soft Hyperedge Selection, referred to as SoftHGNN-SeS, the number of fixed soft hyperedges is set to 16, with an additional 32 dynamic soft hyperedges from which a sparse subset is selected. A sparsity ratio of 50% is applied, meaning that 16 out of the 32 dynamic soft hyperedges are activated during each forward pass. When using SoftHGNN-SeS, the final loss is computed as the sum of load-balancing regularization loss and the original loss of the baseline.
@@
[b] Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param. Gold-YOLO-N [59] 55.7 39.6 12.1 G 2.92 ms 5.6 M Hyper-YOLO-T [15] 54.5 38.5 9.6 G 2.5 ms 3.1 M YOLOv8-N [60] 52.6 37.4 8.7 G 1.77 ms 3.2 M YOLOv9-T [61] 53.1 38.3 7.7 G 2.48 ms 2.0 M YOLOv10-N [62] 53.8 38.5 6.7 G 1.84 ms 2.3 M YOLO11-N [57] 55.3 39.4 6.5 G 1.50 ms 2.6 M YOLO11-N [57] 53.6 38.2 6.5 G 1.59 ms 2.6 M YOLO11-SoftHGNN-N 55.8 ($\uparrow$2.2) 40.2 ($\uparrow$2.0) 8.0 G 1.99 ms 3.2 M YOLOv12-N [58] 55.4 40.1 6.0 G 1.60 ms 2.5 M YOLOv12-N [58] 55.1 39.6 6.0 G 1.63 ms 2.5 M YOLOv12-SoftHGNN-N 57.0 ($\uparrow$1.9) 41.3 ($\uparrow$1.7) 7.4 G 1.99 ms 3.1 M
@@ 5.4.1 Image Classification
In the image classification task, we adopt the classical Vision Transformer (ViT) [4] as our baseline. As illustrated in Figure 2, we incrementally introduce different modules at the end of the ViT encoder for comparison: Self-Attention [4], HGNN [22], the proposed SoftHGNN, and SoftHGNN-SeS (with the Sparse Hyperedge Selection). Table 1 presents the results on the CIFAR-10 [16] and CIFAR-100 [16] datasets. From the table, It can be observed that adding an additional Self-Attention module brings no significant improvement, as ViT itself already relies on attention mechanism. This further highlights the inherent limitation of self-attention in modeling only explicit, pairwise visual associations. When introducing the HGNN module, performance slightly drops on CIFAR-10 and slightly improves on CIFAR-100. This instability is attributed to the hard binary nature of HGNN’s hyperedge assignments. In our experience, HGNNs often require extensive and delicate hyperparameter tuning to achieve competitive performance. In contrast, when the proposed SoftHGNN module is integrated, we observe a significant improvement of 1.9% on CIFAR-10 and 1.8% on CIFAR-100. Moreover, when we increase the number of soft hyperedges and apply sparse hyperedge selection along with the load-balancing regularization, the model achieves further gains. Specifically, SoftHGNN-SeS improves upon the baseline by 2.5% and 2.4% on CIFAR-10 and CIFAR-100, respectively.
@@
Gold-YOLO-S [59] 62.5 45.4 46.0 G 3.82 ms 21.5 M
Hyper-YOLO-S [15] 65.1 48.0 39.0 G 4.7 ms 14.8 M
YOLOv8-S [60] 61.8 45.0 28.6 G 2.33 ms 11.2 M
YOLOv9-S [61] 63.4 46.8 26.4 G 3.46 ms 7.1 M
YOLOv10-S [62] 63.0 46.3 21.6 G 2.49 ms 7.2 M
YOLO11-S [57] 63.9 46.9 21.5 G 2.50 ms 9.4 M
YOLO11-S [57] 62.2 45.6 21.5 G 2.57 ms 9.4 M
@@
YOLOv12-S [58] 64.1 47.6 19.4 G 2.42 ms 9.1 M
YOLOv12-S [58] 63.8 47.0 19.4 G 2.49 ms 9.1 M
@@
Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param.
Gold-YOLO-M [59] 67.0 49.8 87.5 G 6.38 ms 41.3 M
Hyper-YOLO-M [15] 69.0 52.0 103.3 G 9.0 ms 33.3 M
YOLOv8-M [60] 67.2 50.3 78.9 G 5.09 ms 25.9 M
YOLOv9-M [61] 68.1 51.4 76.3 G 6.18 ms 20.0 M
YOLOv10-M [62] 68.1 51.1 59.1 G 4.74 ms 15.4 M
YOLO11-M [57] 68.5 51.5 68.0 G 4.70 ms 20.1 M
YOLO11-M [57] 67.3 50.5 68.0 G 5.04 ms 20.1 M
YOLO11-SoftHGNN-M 68.1 ($\uparrow$0.8) 51.1 ($\uparrow$0.6) 85.0 G 6.07 ms 25.1 M
YOLOv12-M [58] 69.5 52.6 59.8 G 4.27 ms 19.6 M
YOLOv12-M [58] 68.6 51.9 59.8 G 4.75 ms 19.6 M
YOLOv12-SoftHGNN-M 69.2 ($\uparrow$0.6) 52.2 ($\uparrow$0.3) 80.2 G 6.01 ms 26.3 M
5.4.2 Crowd Counting
In the crowd counting task, we adopt the Transformer-based CCTrans [39] as our baseline, and introduce Self-Attention [4], HGNN [22], SoftHGNN, and SoftHGNN-SeS into the multi-scale feature fusion stage, as illustrated in Figure 2. Table 2 presents the experimental results on the ShanghaiTech Part-A and Part-B datasets. We the methods using two different backbone networks: one is the Twins-PCPVT [54] used by the original CCTrans, and the other is the more lightweight and efficient Pyramid Vision Transformer (PVT) [5]. Under both backbones, the proposed SoftHGNN series significantly outperform both Self-Attention and the traditional HGNN. When using Twins-PCPVT as the backbone, SoftHGNN series achieve the largest improvements of 3.6% in MAE and 7.9% in MSE on Part-A, and 7.9% in MAE and 13.1% in MSE on Part-B. When using PVT as the backbone, the improvements brought by the SoftHGNN series are even more substantial: the MAE and MSE improve by up to 3.7% and 7.8% on Part-A, and by up to 10.8% and 16.0% on Part-B. Furthermore, it is worth noting that even with significant improvements, SoftHGNN series still maintains the efficiency of parameters. Specifically, SoftHGNN introduces only about 1M additional parameters, while SoftHGNN-SeS adds approximately 6M. This demonstrates that our proposed SoftHGNN series effectively capture high-order semantic associations that are overlooked by traditional HGNN and Self-Attention, thereby enhancing the model’s representational capacity in an efficient manner.
Figure 3: Detection results on the MS COCO dataset. The input image, ground truth result, and the prediction of YOLO11-N, YOLO11-SoftHGNN-N, YOLOv12-N and YOLOv12-SoftHGNN-N are shown from left to right, respectively.
@@ 5.4.3 Object Detection
For the object detection task, we take the state-of-the-art object detection models YOLO11 [57] and YOLOv12 [58] as baselines. We introduce a SoftHGNN-based feature aggregation and distribution module, named Fuse to SoftHGNN (F2SoftHG), into the neck part of the model, as illustrated in Figure 2. The previous representative real-time detectors Gold-YOLO [59], Hyper-YOLO [15], YOLOv8~10 [60, 61, 62] are also included in the table for comparison.
@@
We conduct experiments on the Nano, Small, and Medium variants of YOLO11 and YOLOv12, as shown in Table 3, Table 4, and Table 5. For the Nano scale, our method improves $\text{AP}_{50}^{val}$ by 2.2% and 1.9%, and $\text{AP}_{50:95}^{val}$ by 2.0% and 1.7% over YOLO11 and YOLOv12, respectively. For the Small scale, it improves $\text{AP}_{50}^{val}$ by 1.6% and 1.0%, and $\text{AP}_{50:95}^{val}$ by 1.5% and 1.0%. For the Medium scale, it achieves an increase of 0.8% and 0.6% in $\text{AP}_{50}^{val}$, and 0.6% and 0.3% in $\text{AP}_{50:95}^{val}$ compared to YOLO11 and YOLOv12, respectively. In addition, our method introduces only a small number of additional parameters and computational cost. Compared with YOLOv12-S, our method introduces only an additional 2.3M parameters and 5.8G FLOPs, while still maintaining real-time performance and high computational efficiency.
@@ 5.5.1 Hyperparameter Sensitivity
Next, we fix the number of heads to 8 and maintain all other settings unchanged, while varying the number of soft hyperedges (4, 8, 16, 32). The results show that SoftHGNN performs better when the number of soft hyperedges is set to 8 or 16, with the best MAE and MSE achieved at 8 soft hyperedges. It is noteworthy that in SoftHGNN, increasing the number of soft hyperedges does not always lead to better performance, despite providing a larger capacity for learnable parameters. We attribute this to two reasons. On the one hand, the number of effective high-order semantic relations is directly related to the complexity of the scenes in the dataset. In most cases, a relatively small number of soft hyperedges (such as 8 or 16) is sufficient to capture enough high-order semantic relations. On the other hand, when the number of soft hyperedges becomes too large, the model tends to capture less important or redundant relations, or overfit to specific bad-case associations. SoftHGNN-SeS is exactly proposed to solve this issue. As shown in Table 1 and Table 2, SoftHGNN-SeS extends the learning capacity of SoftHGNN by first providing a larger number of dynamic candidate soft hyperedges, and then sparsely selecting only the most important ones. This approach allows the model to benefit from a larger hyperedge space while avoiding redundancy and overfitting, thus significantly enhancing the upper bound of its learning ability.

### v3

@@ 5.3 Experimental Setup
[b] Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param. Gold-YOLO-N [goldyolo] 55.7 39.6 12.1 G 2.92 ms 5.6 M Hyper-YOLO-T [hyperyolo] 54.5 38.5 9.6 G 2.5 ms 3.1 M YOLOv8-N [yolov8] 52.6 37.4 8.7 G 1.77 ms 3.2 M YOLOv9-T [yolov9] 53.1 38.3 7.7 G 2.48 ms 2.0 M YOLOv10-N [yolov10] 53.8 38.5 6.7 G 1.84 ms 2.3 M YOLO11-N [yolo11] 55.3 39.4 6.5 G 1.50 ms 2.6 M YOLO11-N [yolo11] 53.6 38.2 6.5 G 1.59 ms 2.6 M YOLO11-SoftHGNN-N 55.8 ($\uparrow$2.2) 40.2 ($\uparrow$2.0) 8.0 G 1.99 ms 3.2 M YOLOv12-N [yolov12] 55.4 40.1 6.0 G 1.60 ms 2.5 M YOLOv12-N [yolov12] 55.1 39.6 6.0 G 1.63 ms 2.5 M YOLOv12-SoftHGNN-N 57.0 ($\uparrow$1.9) 41.3 ($\uparrow$1.7) 7.4 G 1.99 ms 3.1 M
@@ 5.4.1 Image Classification
In the image classification task, we incrementally introduce different modules at the end of the encoder for comparison: Self-Attention [vit], HGNN (construct the hypergraph through $\epsilon$-ball or $k$-NN) [hgnnp], HgVT block [hgvt], the proposed SoftHGNN, and SoftHGNN-SeS (with the Sparse Hyperedge Selection), as illustrated in Figure 2.
On ImageNet-100, we adopt Pyramid Vision Transformer (PVT) [pvt] and Swin-Transformer (SwinT) [swin] as our baseline. As shown in Table 2, since the backbone itself is an Attention architecture, the performance improvement brought by using additional attention modules on a small scale is limited. It only improves the Top-1 accuracy by 0.3% and 1.3% on PVT and SwinT, respectively. When introducing standard HGNN modules, neither the version constructing hypergraphs based on $\epsilon$-ball nor the one based on $k$-NN brings significant improvement. In particular, the $k$-NN-based HGNN decreases the Top-5 performance of PVT by 0.2%. The version introducing HgVT significantly outperforms the former three, achieving performance gains of 2.4% and 1.4% on PVT and SwinT, respectively. However, the version introducing the proposed SoftHGNN has fewer parameters than HgVT but achieves significantly larger performance gains. It achieves Top-1 accuracy improvements of 3.4% and 2.1% on PVT and SwinT, respectively. Finally, the version introducing SoftHGNN-SeS further improves the model’s representational capability by expanding the hyperedge capacity, increasing the Top-1 accuracy of PVT and SwinT by 4.1% and 2.9%, respectively.
On ImageNet-1k, we adopt PVT as the baseline to further verify the applicability of the proposed SoftHGNN on large-scale datasets. As shown in Table 3, the improvement achieved by introducing the proposed SoftHGNN is still significantly higher than other methods, improving the Top-1 and Top-5 accuracy by 0.8% and 0.6% respectively, reaching 75.6% and 92.9%, while introducing only 1.5M parameters. When SoftHGNN-SeS is introduced, the Top-1 and Top-5 accuracy are further improved to 75.9% and 93.1%.
5.4.2 Crowd Counting
In the crowd counting task, we adopt the Transformer-based CCTrans [cctrans] as our baseline, and introduce Self-Attention [vit], HGNN [hgnnp], SoftHGNN, and SoftHGNN-SeS into the multi-scale feature fusion stage, as illustrated in Figure 2. Table 4 presents the experimental results on the ShanghaiTech Part-A and Part-B datasets. We conduct experiments with two different backbone networks: one is the Twins-PCPVT [twins] used by the original CCTrans, and the other is the more lightweight and efficient Pyramid Vision Transformer (PVT) [pvt]. Under both backbones, the proposed SoftHGNN series significantly outperform both Self-Attention and the traditional HGNN. When using Twins-PCPVT as the backbone, SoftHGNN series achieve the largest improvements of 3.6% in MAE and 7.9% in MSE on Part-A, and 7.9% in MAE and 13.1% in MSE on Part-B. When using PVT as the backbone, the improvements brought by the SoftHGNN series are even more substantial: the MAE and MSE improve by up to 3.7% and 7.8% on Part-A, and by up to 10.8% and 16.0% on Part-B. Furthermore, it is worth noting that even with significant improvements, SoftHGNN series still maintains the efficiency of parameters. Specifically, SoftHGNN introduces only about 1M additional parameters, while SoftHGNN-SeS adds approximately 6M. This demonstrates that our proposed SoftHGNN series effectively capture high-order semantic associations that are overlooked by traditional HGNN and Self-Attention, thereby enhancing the model’s representational capacity in an efficient manner.
@@ 5.4.2 Crowd Counting
Gold-YOLO-S [goldyolo] 62.5 45.4 46.0 G 3.82 ms 21.5 M
Hyper-YOLO-S [hyperyolo] 65.1 48.0 39.0 G 4.7 ms 14.8 M
YOLOv8-S [yolov8] 61.8 45.0 28.6 G 2.33 ms 11.2 M
YOLOv9-S [yolov9] 63.4 46.8 26.4 G 3.46 ms 7.1 M
YOLOv10-S [yolov10] 63.0 46.3 21.6 G 2.49 ms 7.2 M
YOLO11-S [yolo11] 63.9 46.9 21.5 G 2.50 ms 9.4 M
YOLO11-S [yolo11] 62.2 45.6 21.5 G 2.57 ms 9.4 M
@@
YOLOv12-S [yolov12] 64.1 47.6 19.4 G 2.42 ms 9.1 M
YOLOv12-S [yolov12] 63.8 47.0 19.4 G 2.49 ms 9.1 M
@@
[b] Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param. Gold-YOLO-M [goldyolo] 67.0 49.8 87.5 G 6.38 ms 41.3 M Hyper-YOLO-M [hyperyolo] 69.0 52.0 103.3 G 9.0 ms 33.3 M YOLOv8-M [yolov8] 67.2 50.3 78.9 G 5.09 ms 25.9 M YOLOv9-M [yolov9] 68.1 51.4 76.3 G 6.18 ms 20.0 M YOLOv10-M [yolov10] 68.1 51.1 59.1 G 4.74 ms 15.4 M YOLO11-M [yolo11] 68.5 51.5 68.0 G 4.70 ms 20.1 M YOLO11-M [yolo11] 67.3 50.5 68.0 G 5.04 ms 20.1 M YOLO11-SoftHGNN-M 68.1 ($\uparrow$0.8) 51.1 ($\uparrow$0.6) 85.0 G 6.07 ms 25.1 M YOLO11-SoftHGNN-M 68.0 ($\uparrow$0.7) 51.0 ($\uparrow$0.5) 81.7 G 5.82 ms 22.8 M YOLOv12-M [yolov12] 69.5 52.6 59.8 G 4.27 ms 19.6 M YOLOv12-M [yolov12] 68.6 51.9 59.8 G 4.75 ms 19.6 M YOLOv12-SoftHGNN-M 69.2 ($\uparrow$0.6) 52.2 ($\uparrow$0.3) 80.2 G 6.01 ms 26.3 M YOLOv12-SoftHGNN-M 69.1 ($\uparrow$0.5) 52.2 ($\uparrow$0.3) 73.6 G 5.09 ms 22.3 M
  • The $\star$ represents that when we input features into the SoftHGNN module, we use $1\times 1$ convolution to compress the channel number and then restore it by using symmetric convolution blocks after the output, in order to reduce the number of parameters.
@@ 5.4.3 Object Detection
For the object detection task, we take the state-of-the-art object detection models YOLO11 [yolo11] and YOLOv12 [yolov12] as baselines. We introduce a SoftHGNN-based feature aggregation and distribution module, named Fuse to SoftHGNN (F2SoftHG), into the neck part of the model, as illustrated in Figure 2. The previous representative real-time detectors Gold-YOLO [goldyolo], Hyper-YOLO [hyperyolo], YOLOv8-v10 [yolov8, yolov9, yolov10] are also included in the table for comparison.
@@
Figure 3: Detection results on the MS COCO dataset. The input image, the prediction of YOLO11-N, YOLO11-SoftHGNN-N, YOLOv12-N, YOLOv12-SoftHGNN-N, and the ground truth are shown from left to right, respectively.
We conduct experiments on the Nano, Small, and Medium variants of YOLO11 and YOLOv12, as shown in Tables 57. For the Nano scale, our method improves $\text{AP}_{50}^{val}$ by 2.2% and 1.9%, and $\text{AP}_{50:95}^{val}$ by 2.0% and 1.7% over YOLO11 and YOLOv12, respectively, while introducing only 0.6M additional parameters and about 1.5G extra FLOPs. For the Small scale, our method improves $\text{AP}_{50}^{val}$ by 1.6% and 1.0%, and $\text{AP}_{50:95}^{val}$ by 1.5% and 1.0%. Meanwhile, it still introduces very little computational overhead. For example, compared with YOLOv12-S, our method introduces an additional 2.3M parameters and 5.8G FLOPs. For the Medium scale, our method achieves an increase of 0.8% and 0.6% in $\text{AP}_{50}^{val}$, and 0.6% and 0.3% in $\text{AP}_{50:95}^{val}$ compared to YOLO11 and YOLOv12, respectively. Since the multi-scale feature maps in Medium models contain substantially more channels than those in Nano and Small models, directly inserting SoftHGNN leads to a higher overhead (about 5M parameters). Thus, we introduce a more parameter-efficient variant (YOLO11-SoftHGNN-M and YOLOv12-SoftHGNN-M) by compressing the channel dimension with a $1\times 1$ convolution before SoftHGNN and restoring it afterward. This lightweight $\star$” variant reduces the overhead (i.e., to only 2.7M additional parameters), while keeping the accuracy essentially unchanged.
@@ 5.5 Ablation Study
5.5.1 Frozen-Encoder Evaluation
To examine whether SoftHGNN can also operate effectively as a lightweight add-on module on top of a fixed image encoder, we conduct an additional frozen-encoder study on ImageNet-100. Following the image classification setting in Section 5.4.1, the compared modules are attached at the end of the encoder. Specifically, we first train the original backbone on ImageNet-100, and then freeze all encoder parameters. After that, we append SoftHGNN and a newly initialized classification head, and optimize only the newly added module and classifier. For comparison, we also report the standard end-to-end training results under random initialization, as well as the original PVT and SwinT results as the baselines.
Table 8: Ablation study on whether to use a frozen backbone trained on ImageNet-100.
Backbone Enc. Frz. Top-1 (%) Top-5 (%)
PVT-Tiny - 82.7 95.4
+SoftHGNN 85.5 96.7
+SoftHGNN 86.1 97.2
SwinT-Tiny - 88.0 97.6
+SoftHGNN 89.5 98.1
+SoftHGNN 90.1 98.3
As shown in Tab. 8, SoftHGNN remains effective even when the image encoder is frozen. On PVT-Tiny, the frozen-encoder setting achieves 85.5% Top-1 and 96.7% Top-5 accuracy, compared with 86.1% and 97.2% under end-to-end training. On SwinT-Tiny, the corresponding results are 89.5% and 98.1%, versus 90.1% and 98.3% for end-to-end training. Although end-to-end optimization still gives the best performance, the degradation in the frozen setting is small, indicating that SoftHGNN can still serve as an effective lightweight adapter without requiring full backbone retraining.
Table 9: Comparison of different SoftHGNN variants on Swin and PVT backbones.
Method Top-1 (%) Top-5 (%)
SwinT-Tiny 88.0 97.6
+SoftHGNN (E-Norm) 90.1 98.3
+SoftHGNN (V-Norm) 89.8 97.9
+SoftHGNN w/o Norm 89.4 97.4
PVT-Tiny 82.7 95.4
+SoftHGNN (E-Norm) 86.1 97.2
+SoftHGNN (V-Norm) 85.7 97.0
+SoftHGNN w/o Norm 85.2 96.7
5.5.2 Ablation on Softmax Normalization
Applying Softmax normalization to the original similarity score matrix $S$ is a crucial step for generating the participation matrix $A$. To investigate the impact of normalization and the two different normalization approaches on the performance of SoftHGNN, we conduct ablation experiments, as shown in Table 9. The results demonstrate that although the model still achieves a substantial improvement over the baseline without Softmax normalization, applying either V-Norm or E-Norm yields significantly better results. This underscores the necessity of the normalization process. Notably, E-Norm slightly outperforms V-Norm; thus, we adopt E-Norm as our default normalization configuration.
@@ 5.5.3 Hyperparameter Sensitivity
Figure 5: The t-SNE visualizations of the feature representation distributions learned by different model variants on ImageNet-100.
Figure 6: Visualization of soft hyperedges on randomly selected validation images for image classification and object detection. The left column displays examples from ImageNet-100 classification, while the right column shows examples from MS COCO object detection.
@@
Next, we fix the number of heads to 8 and maintain all other settings unchanged, while varying the number of soft hyperedges (4, 8, 16, 32). The results show that SoftHGNN performs better when the number of soft hyperedges is set to 8 or 16, with the best MAE and MSE achieved at 8 soft hyperedges. It is noteworthy that in SoftHGNN, increasing the number of soft hyperedges does not always lead to better performance, despite providing a larger capacity for learnable parameters. We attribute this to two reasons. On the one hand, the number of effective high-order semantic relations is directly related to the complexity of the scenes in the dataset. In most cases, a relatively small number of soft hyperedges (such as 8 or 16) is sufficient to capture enough high-order semantic relations. On the other hand, when the number of soft hyperedges becomes too large, the model may introduce overlapping or weakly useful hyperedges, causing semantic diffusion across hyperedges or over-reliance on a few idiosyncratic associations. SoftHGNN-SeS is exactly proposed to address this issue. By first providing a larger pool of candidate hyperedges and then sparsely selecting only the most important ones, together with load-balancing regularization, SoftHGNN-SeS enables the model to benefit from increased hyperedge capacity while encouraging more specialized and balanced hyperedge utilization. As shown in Table 2, 3 and Table 4, SoftHGNN-SeS extends the learning capacity of SoftHGNN by first providing a larger number of dynamic candidate soft hyperedges, and then sparsely selecting only the most important ones. This approach allows the model to benefit from a larger hyperedge space while avoiding redundancy and overfitting, thus significantly enhancing the upper bound of its learning ability.
5.5.4 Qualitative Analysis
We utilize t-SNE to visualize the feature representation distributions learned by different model variants on ImageNet-100, as illustrated in Figure 5. In the visualization, each point represents the feature representation of a sample, with different colors distinguishing different classes. Ten categories are randomly selected for this visualization. Specifically, (1) is the original PVT baseline; (2) is PVT integrated with SoftHGNN, where the number of soft hyperedges is set to 8; (3) also employs SoftHGNN but with the hyperedge count expanded to 48; and (4) is the PVT with SoftHGNN-SeS, with a total of 48 hyperedges (comprising 16 fixed and 32 dynamic hyperedges, with only 16 selected).
Comparing (1) and (2), it is evident that the inclusion of SoftHGNN significantly enhances the separability of feature embeddings across different classes, forming tighter clusters with more distinct boundaries. From (2) to (3), as the number of hyperedges increases substantially, the separability between clusters slightly decreases, which may be attributed to the introduction of irrelevant structures for the task during the expansion of hyperedges. Finally, comparing (3) and (4), the introduction of the SeS strategy leads to significantly superior cluster separability. This suggests that the SeS strategy helps the model focus on task-relevant relational structures, thereby improving the learned feature representations.
Furthermore, we visualize soft hyperedges constructed by SoftHGNN in both image classification and object detection tasks, as illustrated in Figure 6. The left column displays examples from ImageNet-100 classification, while the right column shows examples from MS COCO object detection. In these visualizations, columns $e_{1}$ and $e_{2}$ represent two distinct soft hyperedges. Pixels closer to red indicate higher participation weights in the hyperedge, while pixels toward blue indicate lower participation.
These results clearly demonstrate the behavior of SoftHGNN. For instance, in the second row of the left column, $e_{1}$ primarily models the internal correlations within the bird’s body, whereas $e_{2}$ captures the associations between the bird, surrounding branches, and the background sky. In the first row of the right column, $e_{1}$ models the relationships among the tennis ball, the racket, and the athlete’s body, while $e_{2}$ focuses on the interaction between the athlete and the background environment. In the second row of the right column, $e_{1}$ captures the correlations between the train and the nearby staff, while $e_{2}$ models background elements such as the railway and utility poles. These visualizations indicate that SoftHGNN effectively learns high-order semantic associations within a scene. Moreover, they reveal a form of “division of labor” through semantic differentiation among different hyperedges. This allows the model to capture diverse relational structures, collectively guiding more accurate decision-making for various visual recognition tasks.